OpenStudy (anonymous):

Help? A $700,000 balloon mortgage has terms of 20/4 with a 6% interest rate. What is the monthly payment rounded to the nearest dollar? $4,510 $2,917 $3,092 $5,015

OpenStudy (anonymous):

@ranga

OpenStudy (anonymous):

@agent0smith

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

@BrandonGarza777

OpenStudy (ranga):

This balloon mortgage is amortized over a 20 year period but a huge balloon payment will be due after 4 years. The monthly payment formula is: \[\Large M = \frac{ P * c * (1+c)^{12t}}{ [(1+c)^{12t} - 1] }\]

OpenStudy (ranga):

P is the Principal = $700,000 c is the monthly interest rate in decimal = 6/1200 = 0.005 (Note: I used the symbol c instead of the usual r because normally we are used to using r as an annual interest rate. We should not make the mistake of putting 0.006 in the formula. c is the monthly interest rate in decimal) t is the amortization period term agreed to in the loan. t = 20 years. Plug the numbers into the formula and see what you get for M.

OpenStudy (anonymous):

wait whats c and p?

OpenStudy (anonymous):

ohh hold on

OpenStudy (ranga):

P is Principal or the loan amount. It is $700,000 c is monthly interest rate in decimal. We are given r = 6%. Divide it by 12 to get monthly interest rate in percent 6/12 = 0.5% convert percent to decimal 0.005. So c = 0.005

OpenStudy (anonymous):

I got 37 :(

OpenStudy (ranga):

\[\Large M = \frac{ 700000 * 0.005 * (1.005)^{240} }{ (1.005)^{240} - 1 } = ?\]

OpenStudy (ranga):

The numerator above simplifies to 11585.72 and the denominator computes to 2.31. So M = 11585.72 / 2.31 = ?

OpenStudy (ranga):

$5,015